OpenStudy (anonymous):

Find the mean and the median of the following arithmetic sequences: i. 1, 3, 5, 7, 9,.........,199 ii. 7, 10, 13, 16,....., 607

OpenStudy (tanya123):

Mean = 37.3

OpenStudy (tanya123):

130.6 (for the second)

OpenStudy (anonymous):

The ellipses represents numbers in between. I wasn't sure how to make an equation for this. There has to be a more efficient way to add up the number between 7 and 607, but I'm not sure how to do that.

OpenStudy (tanya123):

Pattern for the second one is 3.

OpenStudy (tanya123):

Keep adding 3 =P

OpenStudy (tanya123):

Till you reach 607 Good Luck!!!

OpenStudy (anonymous):

1,3,5,7,9,....,199 they are in A,P. a=1 common difference d=3-1=2 let n be the number of terms. tn=a+(n-1)d 199=1+(n-1)*2=1+2n-2=2 n-1 2 n=199+1=200 n=200/2=100 mean

OpenStudy (anonymous):

I got 307 for the median for the second problem. Can anyone confirm if this is correct? Also, is the mean 307 as well?

OpenStudy (anonymous):

corect